# Hidden ferromagnetism of centrosymmetric antiferromagnets

### v1

@@ Abstract
The time-reversal symmetry ($\mathcal{T}$) breaking is a signature of ferromagnetism, giving rise to such phenomena as the anomalous Hall effect (AHE) and orbital magnetism (OM). Nevertheless, $\mathcal{T}$ can be also broken in certain classes of antiferromagnets, such as weak ferromagnets or altermagnets, which remain invariant under the spatial inversion. In the light of this similarity with the ferromagnetism, it is tempting to ask whether such anomalous antiferromagnetic (AFM) state can be presented as a simplest ferromagnetic one, i.e. within a minimal unit cell containing only one magnetic cite. We show that such presentation is possible due to the special form of the spin-orbit (SO) interaction in an antiferroelectrically distorted lattice hosting this AFM state. The inversion symmetry, combined with the lattice translations, imposes a severe constraint on the form of the SO interaction, which becomes invariant under the symmetry operation $\{\mathcal{S}|{\bf t}\}$, combining the $180^{\circ}$ rotation of spins ($\mathcal{S}$) with the lattice shift $\bf t$}, connecting antiferromagnetically coupled sublattices. This is the fundamental symmetry property of centrosymmetric antiferromagnets, which justifies the use of the generalized Bloch theorem and transformation to the local coordinate frame with one magnetic cite per cell. It naturally explains the emergence of AHE and OM, and provides transparent expressions for these properties in terms of the electron hoppings and SO interaction operating between nearest neighbors as well as the orthorhombic strain of the next-nearest-neighbor hoppings. The idea is illustrated on a number of examples, using realistic models derived from first-principles calculations. These examples include two-dimensional square lattice, monoclinic VF4 and CuF2, and RuO2-type materials with the tetragonal symmetry.
@@ I Introduction
On the other hand, the AFM materials with broken time-reversal symmetry have been known for decades. Particularly, in 1950s, using purely phenomenological symmetry arguments, Dzyaloshinskii pointed out that there are special types of antiferromagnets, which can host the phenomena of weak ferromagnetism [5], piezomagnetism [6], and magnetoelectricity [7]. Contrary to the regular antiferromagnetism, the magnetic unit cell of these material coincides with the chemical one, being the necessary precondition for macroscopic time-reversal symmetry breaking [1]. Therefore, it is clear that, since there cannot be two independent classes of AFM materials with broken $\mathcal{T}$, the phenomena considered by Dzyaloshinskii should be related to what is now called altermagnetism. In this respect, a very detailed phenomenological classification has been given by Turov, who suggested to divide unconventional antiferromagnets into two major classes, calling them “centrosymmetric” and “centroantisymmetric”, depending on whether the spatial inversion $\mathcal{I}$ enters the magnetic group alone or in the combination with $\mathcal{T}$ [8]. Thus, the centrosymmetric antiferromagnetism in Turov’s classification and encompassing such phenomena as weak ferromagnetism and piezomagnetism has a clear resemblance to what is now called altermagnetism, while the centroantisymmetric antiferromagnetism provides a general framework for understanding the magnetoelectricity. Besides the weak ferromagnetism, piezomagnetism, and magnetoelectricity, Turov consider a wide spectrum of phenomena expected in these unconventional antiferromagnets. Particularly, in 1962, Turov and Shavrov not only predicted the anomalous Hall effect (AHE) in centrosymmetric antiferromagnets, but explicitly stated that it can be driven directly by the AFM order [8, 9]. Turov also pointed out that since magnetoelectricity and weak ferromagnetism belong to two different classes, they are mutually exclusive and do not coexist unless they develop in different magnetic sublattices [10]. The typical example of such coexistence is GdFeO3, where Gd sublattice is magnetoelectric, while Fe sublattice is weakly ferromagnetic [11, 12].
@@
Although many properties of altermagnetic materials were anticipated on the phenomenological level, the recent breakthrough in this field is related to the microscopic understanding of these properties, which became largely possible due to development of first-principles electronic structure calculations. The main attention is focused on the search of the new materials with the large splitting of AFM bands [13, 14, 15, 16, 17, 18, 19, 20, 21, 22], which is regarded as a hallmark of altermagnetism [2, 3, 4, 23].
@@
Such splitting is a consequence of the nonsymmorphic symmetry of crystals, which have several atoms in the chemical cell. The atomic positions in this case are generated by the symmetry operations $\{\mathcal{C}|{\bf t}\}$, combining the rotation $\mathcal{C}$ with the lattice shift $\bf t$}, which connects different sublattices. Then, the AFM order is obtained by combining some of the symmetry operations $\{\mathcal{C}|{\bf t}\}$ with $\mathcal{T}$, so that the magnetic unit cell remains the same as the chemical one, being in line with Dzyaloshinskii’s conjecture [1]. Nevertheless, the new aspect of the problem, which was overlooked in the earlier stages, is that the $\{\mathcal{T}\mathcal{C}|{\bf t}\}$ symmetry gives rise to the spin-splitting of bands in the reciprocal space [2, 3, 23, 24]. This splitting is typically regarded as a signature of ferromagnetism in otherwise AFM materials and believed to be responsible for AHE and other phenomena, which are more typical for ferromagnets.
@@
In our recent work, dealing with the minimal one-orbital model, where the spin-orbit (SO) interaction replicates the form of Dzyaloshinskii-Moriya (DM) interactions in the noncentrosymmetric bonds [5, 25], we have argues that the altermagnetic splitting of bands does not play a key role in AHE and orbital magnetism [26]. Namely, $\mathcal{T}$ can be broken even when the AFM bands are spin-degenerate. Similar behaviour has been found in a more sophisticated multi-orbital model for AFM $\kappa$-type organic conductors and orthorhombically distorted perovskites [17, 19]. Furthermore, we have argued that the fundamental symmetry of such AFM state is $\{\mathcal{S}|{\bf t}\}$, which combines the spin flip $\mathcal{S}=i\hat{\sigma}_{y}$ ($\hat{\sigma}_{y}$ being the Pauli matrix) with the lattice shift $\bf t$} connecting the antiferromagnetically coupled sublattices. This means that the eigenvectors for the spin states $\sigma=\pm$ differ only by a phase factor, which guarantees that (i) the spin bands are degenerate and (ii) the contributions of these bands to the anomalous Hall conductivity are equal to each other and, instead of the partial cancellation, which would occur in regular ferromagnets, we have an addition of such contributions [26].
@@
The time-reversal operation itself is the combination of $\mathcal{S}$ and the complex conjugation $K$, $\mathcal{T}=\mathcal{S}K$. Therefore, if the microscopic Hamiltonian is complex, the operation $\{\mathcal{S}|{\bf t}\}$ is not the same as $\{\mathcal{T}|{\bf t}\}$, which is expected for the regular antiferromagnets. That is why the spin degeneracy can exist even though the time-reversal symmetry is broken. Strictly speaking, the spin degeneracy in this case is not the Kramers’ degeneracy, because the latter implies that the system is invariant under $\mathcal{T}$, which is not the case here.
@@
In this work we will explicitly show that the $\{\mathcal{S}|{\bf t}\}$ symmetry is directly related to the fact that $\mathcal{I}$ is conserved in the centrosymmetric antiferromagnets, which imposes a severe constraint on the form of the SO interaction in magnetic bonds. It is true that $\mathcal{I}$ transforms each sublattice to itself and does not interconnect the sites belonging to different sublattices [27]. However, from the viewpoint of magnetic bonds, $\mathcal{I}$, in the combination with the lattice translations, will unambiguously specify the properties of these bonds over the entire lattice: namely, knowing the bond properties around one magnetic site, one can find the properties around the neighboring site, belonging to different sublattice. Particularly, in the centrosymmetric antiferromagnets, the SO interactions around the sites of different sublattices differ by the sign. Thus, this SO interaction behaves like an AFM object, similar to the Néel field, which leads to the $\{\mathcal{S}|{\bf t}\}$ symmetry of microscopic Hamiltonian.
@@
There is another important aspect of the $\{\mathcal{S}|{\bf t}\}$ symmetry: $\mathcal{S}$ is nothing but the $180^{\circ}$ rotation of spins and this rotation is combined with the shift $\bf t$}, which can be viewed as the lattice translation of a more compact unit cell containing one magnetic site. This constitutes the basis of the so-called generalized Bloch theorem [28], which states that by a unitary transformation to the local coordinate frame, where all spins are pointed in the positive direction of $z$, the AFM system can be described as it would have only one magnetic site in the unit cell, i.e. as a ferromagnet. The generalized Bloch theorem provides a mapping of the AFM system onto a FM one, which naturally explain the breaking of the time-reversal symmetry and all plethora of magnetic effect, which are originally known for the ferromagnets, but can be realized in some AFM systems.
@@
The rest of the article is organized as follows. In Sec. II we will consider the basic symmetry properties of SO interaction imposed by $\mathcal{I}$ in the combination with lattice translations. Then, in Sec. III, we will show that the behavior of the SO interactions is related to even more fundamental properties of the magnetoelectric coupling in antiferroelectrically distorted lattice. Particularly, in order to preserve $\mathcal{I}$, the distortion must be antiferroelectric, which inevitably leads to the doubling of the unit cell and the sign change of the SO interaction when it is considered around two different sublattices. Then, in Sec. IV, we will apply the generalized Bloch theorem and show that, for some components of the SO interaction, the sign change can be compensated by the transformation to the local coordinate frame, so that the AFM system can be formally described as a FM one, with smaller unit cell containing only one magnetic site. In Sec. V, we will introduce the minimal model for the centrosymmetric antiferromagnets with broken $\mathcal{T}$ and derive transparent expressions for the anomalous Hall conductivity and orbital magnetization in the local coordinate frame. Applications for the square lattice, monoclinic VF4 and CuF2, as well as RuO2-type materials with the tetragonal symmetry will be considered in Sec. VI. Finally, in Sec. VII, we will summarize our work. Two appendices deal with the unitary transformation of the SO interaction and elimination of the same-sign components responsible for the weak spin ferromagnetism, and construction of the model Hamiltonian for VF4 and CuF2 on the basis of first-principles electronic structure calculations.
@@ II Basic considerations
The DM interactions are induced by the SO coupling. In the simplest one-orbital model, the SO interactions are given by $\hat{\cal H}^{\rm so}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=i\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot\hat{\boldsymbol{\sigma}}$, where $\hat{\boldsymbol{\sigma}}$ is the vector of Pauli matrices and $\boldsymbol{D}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\sim\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ [29, 30]. Therefore, $\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ and $\boldsymbol{D}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ obey the same symmetry principles.
@@ III Weak Ferromagnetism and Antiferroelectricity
In this section, we will elucidate even more fundamental reason why the DM interactions for the weak ferromagnets have a specific pattern shown in Fig. 1, which preserves the inversion symmetry, but doubles the unit cell, resulting in two sublattices. We believe is that the most fundamental quantity to consider is magnetoelectric coupling $\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$, relating the cross product of spins in the bond to the electric polarization in the same bond, $\vec{P}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot[\boldsymbol{e}_{\boldsymbol{R}}\times\boldsymbol{e}_{\boldsymbol{R}^{\prime}}]$, where $\boldsymbol{e}_{\boldsymbol{R}}$ is the unit vector in the direction of the spin moment at the site $\boldsymbol{R}$ [31, 32, 33].
@@
$\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\equiv[\mathcal{P}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}^{v,c}]$ is the $3\times 3$ tensor. According to our conventions [33], the bold character refers to the spin components ($c$), which couple to the cross product, while the vector symbol refers to the $x$, $y$, and $z$ components ($v$) of the electric polarization. Thus, under the spatial inversion, the $v$ components will transform according to the normal vector rules, while the $c$ components will transform according to the pseudovector rule.
@@
$\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ can be finite even in centrosymmetric bonds. Indeed, the spatial inversion about the center of the bond yields the following property: ${\cal I}\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R}^{\prime},\boldsymbol{R}}=-\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$, which means that $\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ does not necessarily vanish. As the result, the electric polarization $\vec{P}$ can be induced by a noncollinear alignment of spins in otherwise centrosymmetric crystals [31, 32]. Alternatively, electric field $\vec{E}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$, acting in the bond, will interact with $\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ and makes the spins $\boldsymbol{e}_{\boldsymbol{R}}$ and $\boldsymbol{e}_{\boldsymbol{R}^{\prime}}$ noncollinear. The corresponding interaction energy is given by $\boldsymbol{D}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot[\boldsymbol{e}_{\boldsymbol{R}}\times\boldsymbol{e}_{\boldsymbol{R}^{\prime}}]$, where $\boldsymbol{D}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=-\vec{E}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ is the DM interaction induced by the electric field.
@@
In the insulating state, $\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ can be evaluated in terms of the superexchange theory, similar to the DM coupling. Indeed, the latter is given by $\boldsymbol{D}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=\frac{2t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}}{U}$, while the former is $\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=-\frac{2e}{V}\frac{t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\vec{\boldsymbol{r}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}}{U}$, where e-e is the electron charge, $V$ is the unit cell volume, $U$ is the on-site Coulomb repulsion, $\hat{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\hat{\mathbb{1}}+i\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot\hat{\boldsymbol{\sigma}}$ is the transfer integral expanded in terms of the $2\times 2$ unity matrix $\hat{\mathbb{1}}$ and the vector of Pauli matrices, and $\hat{\vec{r}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=\vec{r}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\hat{\mathbb{1}}+i\vec{\boldsymbol{r}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot\hat{\boldsymbol{\sigma}}$ is a similar expansion for the position operator [32].
@@
Next, we consider possible patters of electric fields caused by internal atomic displacements. For simplicity, we assume that all fields are parallel to $z$, but can have different directions in different bonds. For instance, such fields can be due to the buckling of the TM-L-TM bonds, where the intermediate ligand (L) sites are displaced parallel to $z$ relative to the transition-metal (TM) sites. Then, there are three main patterns, which are explained in Fig. 3:

### v2

@@ Abstract
The time-reversal symmetry ($\mathcal{T}$) breaking is a signature of ferromagnetism, giving rise to such phenomena as the anomalous Hall effect (AHE) and orbital magnetism. Nevertheless, $\mathcal{T}$ can be also broken in certain classes of antiferromagnets, such as weak ferromagnets or altermagnets, which remain invariant under the spatial inversion. In the light of this similarity with the ferromagnetism, it is tempting to ask whether such unconventional antiferromagnetic (AFM) state can be presented as a simplest ferromagnetic one, i.e. within the unit cell containing only one magnetic site. We show that such presentation is possible due to special form of the spin-orbit (SO) interaction in an antiferroelectrically distorted lattice hosting this AFM state. The inversion symmetry constrains the form of the SO interaction, which becomes invariant under the symmetry operation $\{\mathcal{S}|{\bf t}\}$, combining the $180^{\circ}$ rotation of spins ($\mathcal{S}$) with the lattice shift $\bf t$}, connecting two antiferromagnetically coupled sublattices. This is the fundamental symmetry property of centrosymmetric antiferromagnets, which justifies the use of the generalized Bloch theorem and transformation to the local coordinate frame with one magnetic site per cell. It naturally explains the emergence of AHE and net orbital magnetization, and provide transparent expressions for these properties in terms of the electron hoppings and SO interaction operating between AFM sublattices, as well as the orthorhombic strain, controlling the piezomagnetic response. The idea is illustrated on a number of examples including two-dimensional square lattice, monoclinic VF4 and CuF2, and RuO2-type materials with the rutile structure, using for these purposes realistic models derived from first-principles calculations.
@@ I Introduction
On the other hand, the AFM materials with broken time-reversal symmetry have been known for decades. Particularly, in 1950s, using phenomenological symmetry arguments, Dzyaloshinskii pointed out that there are special types of antiferromagnets, which can host the phenomena of weak ferromagnetism [5], piezomagnetism [6], and magnetoelectricity [7]. Contrary to the regular antiferromagnetism, the magnetic unit cell of these material coincides with the chemical one, being the necessary precondition for macroscopic time-reversal symmetry breaking [1]. Therefore, the phenomena considered by Dzyaloshinskii should be related to what is now called altermagnetism, as all of them manifest time-reversal symmetry breaking in AFM substances. In this respect, a very detailed phenomenological classification has been given by Turov, who suggested to divide unconventional antiferromagnets into two major classes, calling them “centrosymmetric” and “centroantisymmetric”, depending on whether the spatial inversion $\mathcal{I}$ enters the magnetic group alone or in the combination with $\mathcal{T}$ [8]. Thus, the centrosymmetric antiferromagnetism in Turov’s classification and encompassing such phenomena as weak ferromagnetism and piezomagnetism has a clear resemblance to what is now called altermagnetism, while the centroantisymmetric antiferromagnetism provides a general framework for understanding the magnetoelectricity. Besides the weak ferromagnetism, piezomagnetism, and magnetoelectricity, Turov consider a wide spectrum of phenomena expected in the unconventional antiferromagnets. Particularly, in 1962, Turov and Shavrov have predicted the anomalous Hall effect (AHE) in centrosymmetric antiferromagnets. Moreover, they have explicitly stated that AHE is not a side effect of the weak ferromagnetism. Rather, it is directly driven by the AFM order parameter [8, 9]. Turov also pointed out that since magnetoelectricity and weak ferromagnetism belong to two different classes, they are mutually exclusive and do not coexist unless they develop in different magnetic sublattices [10]. The typical example of such coexistence is GdFeO3, where Gd sublattice is magnetoelectric, while Fe sublattice is weakly ferromagnetic [11, 12].
@@
Although many properties of altermagnetic materials were anticipated on the phenomenological level, the recent breakthrough in this field is related to microscopic understanding of these properties, which became largely possible due to development of first-principles electronic structure calculations. The main attention is focused on the search of the new materials with large splitting of the AFM bands [13, 14, 15, 16, 17, 18, 19, 20, 21, 22], which is regarded as a hallmark of altermagnetism [2, 3, 4, 23].
@@
Such splitting is a consequence of the nonsymmorphic symmetry of crystals with several atoms in cell. The atomic positions in this case are generated by the symmetry operations $\{\mathcal{C}|{\bf t}\}$, which combines a rotation $\mathcal{C}$ with the lattice shift $\bf t$}, which connects different sublattices. Then, the AFM order is obtained by combining some of the symmetry operations $\{\mathcal{C}|{\bf t}\}$ with $\mathcal{T}$, so that the magnetic unit cell remains the same as the chemical one, being in line with Dzyaloshinskii’s conjecture [1]. The new aspect of the problem, which was overlooked in the earlier stages, is that the $\{\mathcal{T}\mathcal{C}|{\bf t}\}$ symmetry gives rise to the spin-splitting of bands in the reciprocal space [2, 3, 23, 24]. This splitting is typically regarded as a signature of ferromagnetism in otherwise AFM materials and believed to be solely responsible for AHE and other phenomena, which are more typical for ferromagnets.
@@
In our recent work, dealing with the minimal one-orbital model, where the spin-orbit (SO) interaction replicates the form of Dzyaloshinskii-Moriya (DM) interactions in the noncentrosymmetric bonds [5, 25], we have argues that the altermagnetic splitting of bands does not play a key role in AHE and orbital magnetism [26]. Namely, $\mathcal{T}$ can be broken even when the AFM bands are spin-degenerate. Similar behaviour has been found in a more sophisticated multi-orbital model for AFM $\kappa$-type organic conductors and orthorhombically distorted perovskites [17, 19]. Furthermore, we have argued that the fundamental symmetry of such AFM state is $\{\mathcal{S}|{\bf t}\}$, which combines the $180^{\circ}$ rotation of spins $\mathcal{S}=i\hat{\sigma}_{y}$ ($\hat{\sigma}_{y}$ being the Pauli matrix) with the lattice shift $\bf t$}, connecting antiferromagnetically coupled sublattices. This means that the eigenvectors for the spin states $\uparrow$ and $\downarrow$ differ only by a phase factor, which guarantees that (i) the spin bands are degenerate and (ii) the contributions of these bands to the anomalous Hall conductivity are equal to each other and, instead of the partial cancellation, which would occur in regular ferromagnets, we have an addition of such contributions [26]. This addition is the fundamentally new effect, which can be of great practical importance.
@@
The time-reversal operation is the combination of $\mathcal{S}$ and the complex conjugation $K$: $\mathcal{T}=\mathcal{S}K$. Therefore, if the microscopic Hamiltonian is complex, the symmetry operation $\{\mathcal{S}|{\bf t}\}$ is not the same as $\{\mathcal{T}|{\bf t}\}$, which is expected for regular antiferromagnets. That is why the spin degeneracy can exist even though the time-reversal symmetry is broken. The spin degeneracy in this case is not the Kramers’ degeneracy, because the latter implies that the system is $\mathcal{T}$-invariant, which is obviously not the case here.
@@
In this work we will explicitly show that the $\{\mathcal{S}|{\bf t}\}$ symmetry is directly related to the fact that $\mathcal{I}$ is conserved in the centrosymmetric antiferromagnets, which imposes a severe constraint on the form of the SO interaction in magnetic bonds. It is true that $\mathcal{I}$ transforms each sublattice to itself and does not interconnect the sites belonging to different sublattices [27]. However, $\mathcal{I}$ will unambiguously specify the properties of magnetic bonds: namely, knowing the bond properties around one magnetic site, one can find the properties around neighboring site, belonging to different sublattice. Particularly, we will explicitly show that the SO interaction on the centrosymmetric antiferroelectric lattice behaves as an AFM object and changes its sign when one moves from one sublattice to another. That is why it collaborates with the AFM Néel field and leads to the $\{\mathcal{S}|{\bf t}\}$ symmetry of microscopic Hamiltonian.
@@
There is another important aspect of the $\{\mathcal{S}|{\bf t}\}$ symmetry: the spin rotation $\mathcal{S}$ is combined with the shift $\bf t$}, which can be viewed as the lattice translation of a more compact unit cell with only one magnetic site. This constitutes the basis of the so-called generalized Bloch theorem [28], which states that by a unitary transformation to the local coordinate frame, where all the spins are pointed in the positive direction of $z$, the AFM system can be described within this compact unit cell as if it would be “ferromagnetic” one. Thus, the generalized Bloch theorem provides a mapping of the AFM system onto a FM one, which naturally explains the breaking of the time-reversal symmetry and emergence of associated with it magnetic effect, which are originally known for ferromagnets, but can be realized in some AFM systems.
@@
The rest of the article is organized as follows. In Sec. II we will consider the basic symmetry properties of SO interaction imposed by $\mathcal{I}$ in the combination with lattice translations. Then, in Sec. III, we will show that the behavior of the SO interactions is related to even more fundamental properties of the magnetoelectric coupling in antiferroelectrically distorted lattice. Particularly, in order to preserve $\mathcal{I}$, the distortion must be antiferroelectric, which inevitably leads to the doubling of the unit cell and the sign change of the SO interaction when it is considered around two sublattices. Then, in Sec. IV, we will apply the generalized Bloch theorem and show that, for some components of the SO interaction, the sign change can be compensated by the transformation to the local coordinate frame, so that the AFM system can be formally described as a FM one, with smaller unit cell containing only one magnetic site. In Sec. V, we will introduce the minimal model for the centrosymmetric antiferromagnets with broken $\mathcal{T}$ and derive transparent expressions for the anomalous Hall conductivity and orbital magnetization in the local coordinate frame. Applications for the square lattice, monoclinic VF4 and CuF2, as well as RuO2-type materials with the tetragonal symmetry will be considered in Sec. VI. In Sec. VII we will explain how different contribution to AHE, associated with spin-degenerate and nondegenerate bands, can be evaluated in realistic electronic structure calculations. Finally, in Sec. VIII, we will summarize our work. Two appendices deal with the unitary transformation of the SO interaction and elimination of the same-sign components responsible for the weak spin ferromagnetism, and construction of the model Hamiltonian for VF4 and CuF2 on the basis of first-principles electronic structure calculations.
@@ II Basic considerations
The DM interactions are induced by the SO coupling. In the simplest one-orbital model, the SO interactions are given by $\hat{\cal H}^{\rm so}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=i\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot\hat{\boldsymbol{\sigma}}$, where $\hat{\boldsymbol{\sigma}}$ is the vector of Pauli matrices and $\boldsymbol{D}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\sim\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ [29, 30]. Therefore, $\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ and $\boldsymbol{D}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ obey the same symmetry rules.
@@ III Weak Ferromagnetism and Antiferroelectricity
In this section, we will elucidate even more fundamental reason why the DM interactions for the weak ferromagnets have a specific pattern shown in Fig. 1, which preserves the inversion symmetry, but doubles the unit cell, resulting in two sublattices. We believe that the most fundamental quantity to consider is magnetoelectric coupling $\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$, relating the cross product of spins in the bond to the electric polarization in the same bond, $\vec{P}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot[\boldsymbol{e}_{\boldsymbol{R}}\times\boldsymbol{e}_{\boldsymbol{R}^{\prime}}]$, where $\boldsymbol{e}_{\boldsymbol{R}}$ is the unit vector in the direction of spin at the site $\boldsymbol{R}$ [31, 32, 33].
@@
$\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\equiv[\mathcal{P}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}^{v,c}]$ is the $3\times 3$ tensor. According to our conventions [33], the bold character refers to the spin components ($c$), which couple to the cross product, while the vector symbol refers to the $x$, $y$, and $z$ components ($v$) of electric polarization. Thus, under the spatial inversion, the $v$ components will transform according to the normal vector rules, while the $c$ components will transform according to the pseudovector rule.
@@
$\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ can be finite in centrosymmetric bonds. Indeed, the spatial inversion about the center of the bond yields the following property: ${\cal I}\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=-\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R}^{\prime},\boldsymbol{R}}=\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$, which means that $\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ does not necessarily vanish. As the result, the electric polarization $\vec{P}$ can be induced by a noncollinear alignment of spins in otherwise centrosymmetric crystals [31, 32]. Alternatively, the electric field $\vec{E}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$, acting in the bond, will interact with $\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ and makes the spins $\boldsymbol{e}_{\boldsymbol{R}}$ and $\boldsymbol{e}_{\boldsymbol{R}^{\prime}}$ noncollinear. The corresponding interaction energy is given by $\boldsymbol{D}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot[\boldsymbol{e}_{\boldsymbol{R}}\times\boldsymbol{e}_{\boldsymbol{R}^{\prime}}]$, where $\boldsymbol{D}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=-\vec{E}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ is the DM interaction induced by the electric field.
@@
In the insulating state, $\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ can be evaluated in terms of the superexchange theory, similar to the DM coupling. Indeed, the latter is given by $\boldsymbol{D}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=\frac{2t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}}{U}$, while the former is $\vec{\boldsymbol{\mathcal{P}}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=-\frac{2e}{V}\frac{t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\vec{\boldsymbol{r}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}}{U}$, where e-e is the electron charge, $V$ is the unit cell volume, $U$ is the on-site Coulomb repulsion, $\hat{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\hat{\mathbb{1}}+i\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot\hat{\boldsymbol{\sigma}}$ is the transfer integral expanded in terms of the $2\times 2$ unity matrix $\hat{\mathbb{1}}$ and the vector of Pauli matrices, and $\hat{\vec{r}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=\vec{r}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\hat{\mathbb{1}}+i\vec{\boldsymbol{r}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot\hat{\boldsymbol{\sigma}}$ is similar expansion for the position operator [32].
@@
Next, let us consider possible patters of electric fields caused by internal atomic displacements. For simplicity, we assume that all fields are parallel to $z$, but can have different directions in different bonds. For instance, such fields can be due to the buckling of the TM-L-TM bonds, where the intermediate ligand (L) sites are displaced parallel to $z$ relative to the transition-metal (TM) sites. Then, there are three main patterns, which are explained in Fig. 3: